OpenStudy (anonymous):

Solve the system by substitution. {2x-y+z=-4 {z=5 {-2x+3y-z=-10

OpenStudy (kamille):

So, you have that z=5. Now change everywhere, where is z, to 5. what do you get?

OpenStudy (anonymous):

2x-y+5=-4 -2x+3y-5=-10

OpenStudy (kamille):

so, now you should substitute one variable with another. It should be done leaving one variable (without any numbers or any minuses near it): 2x-y+5=-4 -y=-4-5-2x -y=-9-2x|:(-1) (I do this to eliminate the minus near variable "y". Don't forget, when you divide a number by a negative number, it changes it sign) y=9+2x No, insert 9+2x into the third equation and solve it. What do you get?

OpenStudy (anonymous):

-2x+3(9+2x)-5=-10 like this?

OpenStudy (kamille):

yes, you are absolutely right. NDo you know how to solve this equation?

OpenStudy (anonymous):

not really

OpenStudy (kamille):

okay, I will help you. -2x+3(9+2x)-5=-10 -2x+27+6x-5=-10 -2x+6x+22=-10 4x=-32|:4 x=-8 Do you understand how I did it?

OpenStudy (anonymous):

I think so!! what would be the equation for the first problem?

OpenStudy (kamille):

um,sorry, I didn't uderstand? So, now you have that x=-8 and z=5, right? Now you have to find the y... Don't forget, that you have equation y=9+2x so y is: y=9+2*(-8)=9-16=-7 Your answer is (-8;-7;5)

OpenStudy (anonymous):

okay i get it thank you so much!

OpenStudy (kamille):

you are welcome;))